# Berms without Calibration

### v1

@@ 5 Stochastic Ratios of Annuities
To reflect the dependence of a Bermudan swaption price on stochastic ratios of annuities we suggest using Theorem 2 from Section 1. The convexity adjustments of the long and short swap rates will translate into the corresponding convexity adjustment of the forward starting swap rate (see the conversion formulas in [5]). Therefore, a single shift parameter per each $X_{i}$ from the previous section will suffice to control the stochastic ratios at least in the first order.
@@ Conclusion
We developed an alternative model for Bermudan swaption valuation which allows us to express explicitly the dependence of Bermudan swaption price on the swap rates distributions and correlations between them. The model replaces the use of unobservable mean reversion parameters as in traditional approaches by swap rates correlations which can be estimated from historical data or implied from midcurve swaptions and CMS spreads markets. The model does not require product specific calibrations, and its lattice style analogy has only that many steps as the number of swaption exercises. The approach offers potential to a faster and more accurate risk management of the Bermudan swaptions.
One could try to trade directly off this model, but they would need to maintain adjustments between three separate markets: bermudans, midcurves and swaptions with relative strikes. This is because all the three markets are exotic and supply/demand factors play significant role in addition to no arbitrage principle.
Acknowledgement. The author would like to express a sincere gratitude to Andrew Green, Shuqing Ma and Scotiabank Innovation Hub for supporting this research. I am also grateful to various Scotiabank GAFE team members for creating friendly research environment and for useful discussion with Chris Dennis, Bin Guo, Michael Konikov, Dirk Petera, Duan Qian and Alex Zemlianov.

### v2

@@ 5 Stochastic Ratios of Annuities
While Theorem 8 and its derivation do not depend on the deterministic annuity ratio assumption, the numerical implementation does require to choose the dynamics for the ratios. For the numerical experiment in the previous section we assumed that the annuity ratios are deterministic and the Hagan lattice style model showed a good fit to the market.
The deterministic annuity ratio assumption is widely used in financial engineering (see, for example, [3]). It relies on empirical observation that an annuity ratio has a low variance and is an approximately slow changing linear function of the corresponding long and short swap rates. Nevertheless, effects like mild correlation skews in midcurve swaptions and CMS spread options markets make it beneficial to have a handle on the stochastic part of annuities ratios.
For a Bermudan swaption to reflect the dependence of the price on stochastic ratios of annuities as an alternative to the correlation skew management we suggest using Theorem 2 from Section 1. The convexity adjustments of the long and short swap rates will translate into the corresponding convexity adjustment of the forward starting swap rate (see the conversion formulas in [6]). Therefore, a single distribution shift/convexity adjustment parameter per each coterminal swap rate distribution would suffice to control the stochastic ratios at least in the first order.
@@ Conclusion
We developed an alternative model for Bermudan swaption valuations which allows us to express explicitly the dependence of Bermudan swaption price on the swap rate distributions and correlations between them. The model replaces the use of mean reversion parameters (as in traditional approaches) with swap rate correlations, which can be estimated from historical data or implied from midcurve swaption and CMS spread markets. The model does not require product specific calibrations, and its lattice style analogy has only that many steps as the number of swaption exercises. The approach offers potential to a faster and more accurate risk management of Bermudan swaptions. We provided a numerical implementation based on the implied volatility-by-strike swap rate distribution parameterisation and the deterministic annuity ratio assumption. The model demonstrated a good fit to the market.
Acknowledgement. The author would like to express a sincere gratitude to Andrew Green, Shuqing Ma and Scotiabank Innovation Hub for supporting this research. I am also grateful to various Scotiabank GAFE team members for creating a friendly research environment and for useful discussion with Chris Dennis, Bin Guo, Michael Konikov, Dirk Petera, Duan Qian, Rex Sutton and Alex Zemlianov.